• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name averett

    Log entry time 00:14:57 on July17,2003

    Entry number 104154

    keyword=Swing Shift Summary



    16:00 Access in progress to install BeO target.
    16:30 Access ended.
    17:00 MCC doing beam dump ion chamber studies.
    19:00 MCC requests we ramp Q1R. They will allow us to run up to 5uA tonight.
    Begin ramping Q1R to1276.88Amps.
    19:25 MCC tests done. Begin ramping septum. Perform Bullseye scan with 2 uA beam. Single carbon foil.
    19:45 Begin Bullseye scan

    HARP run DAQ run BPMB
    2133 21362/2327 -1.2, 0.0
    2134 21363/2328 -3.0, 0.0
    2135 21364/2329 -3.0, -2.0
    2136 21365/2330 -3.0, 2.0
    2137 21366/2331 -1.0, 2.0
    2138 21367/2332 -1.0, -2.0
    2139 21368/2333 1.0, -2.1
    2140 21639/2334 1.1, 0.0
    2141 21640/2335 1.0, 2.0
    2142 21641/2336 -1.2, 0.1

    20:35 Finish bullseye scan. MCC measures bleedthrough current from Halls B
    and C. Bleedthrough =19%.
    20:45 Elastic run 21373/2338
    21:00 Lower right arm momentum to 1.9224 GeV, septum to I=178A.
    Run 21374/2339
    21:45 Request access to remove BeO. Perform NMR/EPR. Set HRS-right to P=2.093 GeV for QE kinematics. Septum I=193 Amps.
    22:30 Access finished, waiting for beam restore.
    22:50 Beam restored, begin QE runs.